ADNAN MUHAMMED

Grade 12,

If Z1 and Z2 are two complex nos. such that Z1+Z2i=Z2 THEN origin, Z1 and Z2 r the vertices of a___________________ TRIANGLE?

If Z1 and Z2 are two complex nos. such that Z1+Z2i=Z2 THEN origin, Z1 and Z2 r the vertices of a___________________ TRIANGLE?

Grade:

2 Answers

mycroft holmes
13 years ago

The lengths of the three sides of the triangle are 

 

We see that 

 

Also, 

 

This also tells us that 

 

Hence the triangle is a right angled isoceles triangle

Priyansh Bajaj AskiitiansExpert-IITD
13 years ago

Dear Prashant,

Solution:- z1 + z2 i = z2

          or, z2 - z1  = i z2

          or, (z2 - z1) / (z2 - 0) = i ['0' is written to represent origin]

or, (z2 - z1) / (z2 - 0) = e(i∏/2) [|z2 - z1| / |z2 - 0|]; where,|z2 - z1| = |z2 - 0| and  

                    therefore,[|z2 - z1|/|z2 - 0|] =1

 

Hence, origin, z1 and z2 are the vertices of isosceles right angled triangle, having right angle between (z2 - z1) and (z2 - origin) sides.


Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

All the best Prashant!!!

Regards,
Askiitians Experts
Priyansh Bajaj